### v2

# PocketGS: On-Device Training of 3D Gaussian Splatting for High Perceptual Modeling

Abstract
Efficient and high-fidelity 3D scene modeling is a long-standing pursuit in computer graphics. While recent 3D Gaussian Splatting (3DGS) methods achieve impressive real-time modeling performance, they rely on resource-unconstrained training assumptions that fail on mobile devices, which are limited by minute-scale training budgets and hardware-available peak-memory. We present PocketGS, a mobile scene modeling paradigm that enables on-device 3DGS training under these tightly coupled constraints while preserving high perceptual fidelity. Our method resolves the fundamental contradictions of standard 3DGS through three co-designed operators: $\mathcal{G}$ builds geometry-faithful point-cloud priors; $\mathcal{I}$ injects local surface statistics to seed anisotropic Gaussians, thereby reducing early conditioning gaps; and $\mathcal{T}$ unrolls alpha compositing with cached intermediates and index-mapped gradient scattering for stable mobile backpropagation. Collectively, these operators satisfy the competing requirements of training efficiency, memory compactness, and modeling fidelity. Extensive experiments demonstrate that PocketGS is able to outperform the powerful mainstream workstation 3DGS baseline to deliver high-quality reconstructions, enabling a fully on-device, practical capture-to-rendering workflow.
Figure 1: Our PocketGS enables high-quality end-to-end 3DGS reconstruction on commodity smartphones. Compared to standard 3DGS workstation baselines, PocketGS achieves superior visual fidelity (LPIPS: 0.108) within a tight training budget ( 500 iterations, $\sim$4 minutes on an iPhone 15).
I Introduction
3D Gaussian Splatting (3DGS) [20] is a promising paradigm for high-quality scene modeling, advancing mixed reality [15, 22], digital twins [19, 39], and robotic simulation [7, 40]. By replacing implicit neural representations [24] with rasterization-friendly 3D Gaussians, it achieves real-time high-fidelity modeling, but still suffers from high computation, memory overhead, and training time, hindering on-device training for mobile 3D content creation.
Training 3DGS on device introduces substantial challenges, as it must deliver high perceptual quality under the limited computational resources and strict thermal constraints of mobile platforms. Consequently, the training process must meet on-device constraints: 1) completion within a few minutes ( $\textless 5$ min), 2) respecting the hardware-available memory budget ( $\textless 3$ GB).
These constraints require rethinking 3DGS under resource-limited regimes beyond conventional resource-unconstrained assumptions. Gaussian-based modeling typically includes: (1) recovering camera parameters and a point cloud from images, (2) initializing 3D Gaussians from this geometry, and (3) optimizing the Gaussians via differentiable rendering for scene recovery.
Existing designs assume ample time and memory across all stages, but on-device settings break these assumptions, yielding three tightly coupled contradictions.
Input-Recovery Contradiction. A primary limitation of on-device 3DGS arises from the contradiction between unreliable geometric input and accurate scene recovery. Mobile scans provide coarse geometry as imprecise poses and sparse points, causing noisy initialization. Existing methods compensate by densifying Gaussians, but this raises computation and memory overhead, leading to long training times and memory overflow. Consequently, on-device 3DGS requires reliable geometric input to minimize reliance on excessive densification for scene recovery.
Initialization-Convergence Contradiction. Another key limitation lies in the contradiction between heuristic initialization and stable convergence. Existing methods disregard geometric priors and initialize input points as isotropic 3D Gaussians, relying on prolonged optimization to compensate for the resulting deficiencies. Under on-device constraints, extra iterations are costly and violate time, energy, and thermal budgets. Thus, on-device 3DGS requires geometry-conditioned initialization to reduce optimization burden, thereby facilitating stable convergence.
Hardware-Differentiability Contradiction. On-device 3DGS is at odds with mobile execution: differentiable rendering requires storing intermediates for backpropagation, but mobile pipelines restrict gradient access through opaque blending, suffer from costly GPU–CPU synchronization in optimization, and incur heavy data movement and state misalignment from depth-ordered alpha blending. Therefore, on-device 3DGS requires a mobile-oriented differentiable rendering framework that bounds and replays only the necessary intermediates for Gaussian optimization.
In this work, we present PocketGS, a 3DGS training paradigm that achieves high-perceptual modeling under on-device constraints via three co-designed operators (Fig. LABEL:top,2). To address the Input-Recovery Contradiction, the geometry-prior construction operator $\mathcal{G}$ builds a geometry-faithful prior, providing the required geometric guidance without costly in-loop Gaussian densification beyond the on-device envelope. To resolve the Initialization-Convergence Contradiction, the prior-conditioned Gaussian parameterization operator $\mathcal{I}$ uses this prior to shape initial Gaussians, improving conditioning for stable, iteration-limited optimization and reducing degradation from noise-sensitive initialization. Finally, to overcome the Hardware-Differentiability Contradiction, the hardware-aligned splatting optimization operator $\mathcal{T}$ makes differentiable rendering practical on mobile GPUs, preserving correct gradients and stable updates without prohibitive memory traffic or CPU–GPU synchronization. Extensive experiments show that PocketGS achieves perceptual quality superior to mainstream workstation 3DGS baselines, while operating within on-device constraints.
Our main contributions are:
  • We introduce a geometry-prior construction operator $\mathcal{G}$ that provides compact yet reliable geometric inputs for on-device 3DGS, reducing the dependence on heavyweight reconstruction and costly training-time densification.
  • We propose a prior-conditioned Gaussian parameterization operator $\mathcal{I}$ that improves initialization quality and optimization conditioning, enabling stable convergence under a strict on-device iteration budget.
  • We develop a hardware-aligned splatting optimization operator $\mathcal{T}$ that renders differentiable splatting practical on mobile GPUs, supporting accurate gradients and stable updates within hardware memory and bandwidth constraints.
  • We validate that, while satisfying on-device constraints, PocketGS achieves competitive perceptual quality relative to standard workstation 3DGS pipelines on publicly available benchmarks and our self-collected MobileScan datasets.
Figure 2: Overview of the PocketGS framework. PocketGS tackles on-device 3DGS training through three coupled operators: ($\mathcal{G}$) Geometry Prior Construction employs an information-gated gate for frame selection, followed by GPU-native Schur-complement BA and single-reference MVS. The MVS module constructs a 3D cost volume by sampling depth hypotheses $d_{1},\dots,d_{n}$} via census matching to produce a dense geometric scaffold. ($\mathcal{I}$) Prior-Conditioned Parameterization seeds anisotropic Gaussians by estimating local surface statistics (normals $\mathbf{s}_{n}$ and scales) to front-load structure discovery via disc-like covariance seeding. ($\mathcal{T}$) Hardware-Aligned Splatting implements a mobile-native differentiable renderer using unrolled alpha-compositing ($S={C_{in},\alpha}$) and index-mapped gradient scattering to ensure stable backpropagation within tight mobile memory bounds of the canonical parameter buffer ($\mu,\Sigma,c,\alpha$).
II Related Work
II-A Novel View Synthesis
Novel view synthesis targets photorealistic rendering from unseen viewpoints [33, 16]. The field has progressed from implicit NeRF-style radiance fields [24, 3, 1] to more efficient explicit voxel/hash-grid representations [13, 32, 25, 2, 17], and further to 3DGS with rasterization-friendly Gaussians for fast training and rendering [20, 8, 9]. Despite advances [27, 36, 35, 19, 4, 18, 5, 10], most 3DGS pipelines still rely on offline SfM [34, 31] for pose and sparse geometry, whose high cost and incomplete coverage hinder practical on-device training in resource-constrained settings.
II-B Mobile 3D Scanning
Mobile 3D scanning captures images on mobile devices and reconstructs real-world 3D structure, typically via visual-inertial odometry for pose estimation and multi-view stereo for geometry [28, 29, 14, 38]. Systems such as ARKit and ARCore follow this pipeline [21, 11, 26] but are limited by point/mesh representations, whose sparse and noisy reconstructions lack the capacity for high-fidelity rendering [33, 16]. Meanwhile, novel view synthesis has shown superior modeling and rendering quality, opening new opportunities for mobile 3D scanning [33, 20]; however, existing efforts largely emphasize mobile rendering efficiency [6, 30], leaving on-device training underexplored.
III The PocketGS Paradigm
PocketGS enables on-device 3DGS training: given a short mobile capture $I_{t}$}, it optimizes a Gaussian scene $\Theta$ under strict runtime and peak-memory budgets on tile-based mobile GPUs. It co-designs three coupled operators: $\mathcal{G}$ derives a compact geometry prior $\mathcal{P}$ from $I_{t}$}, $\mathcal{I}$ uses $\mathcal{P}$ to initialize well-conditioned Gaussians $\Theta_{0}$, and $\mathcal{T}$ performs hardware-aligned differentiable splatting to keep training GPU-resident with correct gradients. Formally, $\mathcal{G}:(\{I_{t}\},\{\hat{T}_{t}\})\mapsto(\{T_{t}\},\mathcal{P}),\ \mathcal{I}:\mathcal{P}\mapsto\Theta_{0},$ and $\mathcal{T}:(\Theta_{0},\{I_{t},T_{t}\})\mapsto\Theta^{*}$, balancing perceptual quality with on-device constraints.
III-A Geometry Prior Construction under Resource Budgets
To resolve the Input-Recovery Contradiction, Operator $\mathcal{G}$ constructs a low-noise and low-redundant 3DGS-aligned prior within mobile compute and memory constraints. Treating BA (Bundle Adjustment) and MVS (Multi-View Stereo) as core prior-construction primitives, their outputs directly condition the 3DGS optimization, mitigating over-reliance on excessive in-loop Gaussian densification.
III-A1 Information-Gated Frame Subsampling
To reduce input noise and provide high-quality photometric ground truth for 3DGS training, PocketGS adopts a real-time three-stage gating mechanism that admits only frames with incremental geometric utility, maximizing information yield per compute unit.
Displacement Gate. We ensure sufficient parallax for robust triangulation by computing the camera displacement $d$ between the current frame and the last selected keyframe, $d=\|\mathbf{t}_{curr}-\mathbf{t}_{last}\|_{2}$. A frame is considered only if $d\geq\tau_{d}$, where $\tau_{d}=0.05m$, suppressing redundant viewpoints while maintaining tracking stability.
Sharpness Heuristic. We reject frames degraded by motion or defocus blur using an efficient approximation of gradient energy. We estimate the sharpness score $S$ by summing absolute luma differences over a sparsely sampled grid $\Omega$:
$S=\frac{1}{|\Omega|}\sum_{(x,y)\in\Omega}\left(|I(x+\Delta,y)-I(x,y)|+|I(x,y+\Delta)-I(x,y)|\right)$ (1)
where $\Delta$ is the step size. This score correlates with high-frequency content and favors sharp frames.
Candidate Windowing. We select the locally optimal frame from a short temporal window (8 frames or 250ms). A candidate replaces the current best only if $S_{new}>(1+r)S_{best}$, with $r=0.05$. This commits the sharpest representative per viewpoint and bounds downstream optimization workload.
Keyframe admission is explicitly bounded by the three gates, reducing downstream BA/MVS workload roughly proportional to admitted keyframes rather than raw capture length, thereby constraining both runtime and peak intermediate storage.
III-A2 GPU-Native Global BA as Mobile MAP Refinement
To address mobile input imprecision, PocketGS implements GPU-native global BA to jointly optimize poses and 3D points, delivering 3DGS-required precision without over-densification. The robustified reprojection objective for PocketGS:
$\min_{\{\mathbf{T}_{i}\},\{\mathbf{P}_{j}\}}\sum_{i,j}\rho\left(\|\pi(\mathbf{T}_{i},\mathbf{P}_{j})-\mathbf{p}_{ij}\|^{2}_{\Sigma_{ij}}\right)$ (2)
where $\pi$ is the projection function, $\mathbf{T}_{i}\in SE(3)$ is the pose of camera $i$, $\mathbf{P}_{j}\in\mathbb{R}^{3}$ is the 3D position of point $j$, $\mathbf{p}_{ij}\in\mathbb{R}^{2}$ is the observed 2D projection, and $\rho$ is the robust Huber loss.
Scale-Aware Gauge Fixing. To resolve the 7-DoF ambiguity of monocular reconstruction, we fix the 6-DoF pose of the first keyframe and fix scale by constraining the translational component of the second keyframe along the dominant baseline axis, yielding a consistent metric frame.
GPU-Native Schur Complement. The BA bottleneck is solving the normal equations $\mathbf{H}\Delta=\mathbf{b}$, where $\mathbf{H}=\mathbf{J}^{T}\mathbf{J}$. We exploit the sparse block structure via a full GPU Schur complement. Partitioning parameters into camera pose (c) and point (p) blocks yields the reduced camera system:
$(\mathbf{H}_{cc}-\mathbf{H}_{cp}\mathbf{H}_{pp}^{-1}\mathbf{H}_{pc})\Delta_{c}=\mathbf{b}_{c}-\mathbf{H}_{cp}\mathbf{H}_{pp}^{-1}\mathbf{b}_{p}$ (3)
Since $\mathbf{H}_{pp}$ is block-diagonal with independent $3\times 3$ blocks per point, $\mathbf{H}_{pp}^{-1}$ is computed in parallel across points on the GPU. After solving for $\Delta_{c}$, we recover $\Delta_{p}$ by back-substitution.
Iterative Geometric Refinement. BA is embedded in an iterative refinement loop. After each global step, we re-triangulate features using updated poses, filter points with high reprojection error or insufficient triangulation angles, and remove observations behind the camera. This purification reduces outliers and stabilizes dense reconstruction.
Schur complement reduces solving from the full joint system to a smaller camera system, while $\mathbf{H}_{pp}^{-1}$ is computed in parallel; the design keeps memory dominated by sparse observations rather than dense matrices, making global refinement feasible on-device.
III-A3 Single-Reference Cost-Volume MVS
To address mobile input sparsity, PocketGS’s lightweight MVS generates a dense geometry-faithful point cloud, providing structural richness for Gaussian initialization without over-densification.
Probabilistic Depth Range Estimation. Plane-sweep efficiency depends on depth range. For each target frame, we project sparse BA points and compute the 5% and 95% depth quantiles $q_{.05}$ and $q_{.95}$. We define the search range as $q_{.05}\cdot 0.6,q_{.95}\cdot 1.6]$, focusing computation on the most probable volume.
Memory-Efficient Cost Volume. To minimize memory footprint, we construct the cost volume using only a single, optimally chosen reference frame. The reference maximizes
$S_{ref}=\exp\left(-\frac{(b-b_{target})^{2}}{2\sigma_{b}^{2}}\right)\cdot\max\left(\frac{\alpha}{\alpha_{min}},1\right)$ (4)
balancing baseline length $b$ against a target baseline $b_{target}$ and enforcing sufficient viewing angle $\alpha$. The matching cost is computed using the Census Transform [12] for robustness to illumination changes. Depth is inferred via plane sweep [37] with Semi-Global Matching aggregation. We filter depth maps using a confidence threshold of 0.4 and fuse them into a dense point cloud, which serves as the geometric scaffold for $\mathcal{I}$.
III-B Prior-Conditioned Gaussian Parameterization
To ensure stable convergence under a strict mobile iteration budget, operator $\mathcal{I}$ improves early-stage conditioning by replacing isotropic initialization with prior-conditioned, surface-aligned anisotropic Gaussians, thereby injecting structural cues from $\mathcal{G}$ and reducing reliance on photometric signals to infer geometry.
III-B1 Local Surface Statistics Estimation
To inject geometric prior knowledge into initialization, we estimate local surface statistics for each point $\mathbf{p}_{i}$ in the dense MVS point cloud via KNN (K=16). The neighborhood covariance matrix is computed as:
$\mathbf{C}=\frac{1}{K}\sum_{k=1}^{K}(\mathbf{p}_{k}-\bar{\mathbf{p}})(\mathbf{p}_{k}-\bar{\mathbf{p}})^{T}$ (5)
where $\bar{\mathbf{p}}_{i}$ is the centroid of the neighborhood.The eigenvector corresponding to $\mathbf{C}_{i}$’s smallest eigenvalue robustly estimates the local surface normal $\mathbf{n}_{i}$. This GPU-parallelized process provides critical structural cues for informed initialization.
Fixed local neighborhood size ensures embarrassingly parallel computation with bounded per-point memory, aligning with PocketGS’s mobile constraints.
III-B2 Disc-Like Covariance Seeding
We initialize each Gaussian as a thin, disc-like ellipsoid tangent to the estimated surface, directly addressing the contradiction by replacing naive isotropic initialization with geometry-conditioned anisotropy. The tangential scale $s_{t}$ is derived from local point density, estimated as the average distance to the K=3 nearest neighbors:
$s_{t}=\frac{1}{3}\sum_{k=1}^{3}\|\mathbf{p}-\mathbf{p}_{k}\|$ (6)
We set the normal-direction scale $s_{n}$ as a fraction of tangential scale:
$s_{n}=s_{t}\times r_{normal}$ (7)
where $r_{normal}=0.3$. Scale parameters are optimized in log-space, $\mathbf{s}_{log}=\log(\mathbf{s})$, ensuring positivity and numerical stability. The initial rotation $\mathbf{q}$ aligns the Gaussian local z-axis with the estimated surface normal $\mathbf{n}_{i}$. Opacity is initialized to a fixed low value (logit of 0.1). This geometry-conditioned anisotropy supplies strong priors on surface orientation and extent, accelerating convergence under limited iterations.
III-C Hardware-Aligned Differentiable Splatting Optimization
Operator $\mathcal{T}$ rewrites the 3DGS training loop for GPU-resident optimization on mobile GPUs under tight memory/bandwidth budgets, reducing intermediate storage and synchronization overhead while enabling differentiable blending. By unrolling compositing and decoupling parameter/state layout from rendering order, it supports end-to-end on-device training with correct, stable gradients.
III-C1 Unrolled Alpha-Compositing with Forward Replay Cache
To address hidden intermediate states, PocketGS unrolls alpha-compositing into an explicit differentiable pipeline with cached intermediates. Standard front-to-back compositing follows:
$C_{out}=C_{in}(1-\alpha)+\alpha c,$ (8)
where $C_{in}$ is the pre-composite color, and $(\alpha,c)$ are the opacity and color of the current Gaussian fragment. In a mobile rendering pipeline, hardware blending returns only $C_{out}$ while hiding intermediate factors, making correct gradient propagation unavailable without costly framebuffer reads.
Explicit Computation Graph.
We process visible Gaussians in depth-sorted batches. Each batch explicitly takes the previous accumulator as input, exposing the intermediates necessary for the backward pass:
$\frac{\partial\mathcal{L}}{\partial C_{in}}=\frac{\partial\mathcal{L}}{\partial C_{out}}(1-\alpha),\quad\frac{\partial\mathcal{L}}{\partial\alpha}=\frac{\partial\mathcal{L}}{\partial C_{out}}(c-C_{in}).$ (9)
Bandwidth-Efficient Cache.
To minimize memory bandwidth pressure-a critical constraint for mobile energy efficiency-we introduce a forward replay cache. Instead of caching full per-pixel splat lists, we store a minimal set $S=\{C_{in},\alpha\}$ and use a per-pixel counter buffer. By resetting only the $O(WH)$ counter rather than the entire $O(WHK_{max})$ cache, we significantly reduce the bandwidth overhead during the backward traversal.
III-C2 Index-Mapped Gradient Scattering
Maintaining optimizer consistency during depth-sorted rendering typically requires physical memory reordering, which incurs heavy data movement. We resolve this by decoupling the sorted view from the canonical layout. A GPU-based sort generates an index list $\pi$, allowing the renderer to access parameters $\theta_{\pi(i)}$ while they remain stationary in memory. During backpropagation, gradients are scattered back to the canonical buffers:
$\nabla\theta_{\pi(i)}+=g_{i},$ (10)
ensuring that Adam optimizer states remain aligned with their respective parameters without CPU intervention or redundant memory copies.
III-C3 On-GPU Parameter Optimization
To eliminate the synchronization latency of CPU-driven updates, the entire optimization loop is executed on-device. We maintain Adam states (moments $m,v$) in dedicated GPU buffers and perform updates via a unified compute kernel. To ensure convergence under limited FP16 precision, we employ specialized parameterization: Opacity uses logit-space updates with a micro-step threshold; Scale is optimized in log-space with explicit clamping; and Rotation is updated via tangent-space projection and sign-unification. This fully GPU-resident schedule minimizes the host-device communication footprint while maximizing hardware utilization.
IV Experiments
Figure 3: Qualitative comparison on our MobileScan dataset. PocketGS (Ours) consistently recovers sharper textures and finer details compared to baselines. Notably, 3DGS-MVS-WK suffers from artifacts despite its high-density prior, as the over-concentrated initial points impede effective Gaussian redistribution within a limited iteration budget. Our method avoids this bottleneck through prior-conditioned initialization, achieving superior structural fidelity (e.g., bicycle spokes) that closely matches the ground truth (GT).
LLFF NeRF-Synthetic MobileScan
[…]
3DGS-SFM-WK 21.01 0.641 0.405 108.0 18k 21.75 0.800 0.243 83.7 12k 21.16 0.687 0.398 112.8 23k
3DGS-MVS-WK 19.53 0.637 0.387 313.1 40k 24.47 0.887 0.128 532.1 50k 20.85 0.781 0.281 534.5 165k
PocketGS (Ours) 23.54 0.791 0.222 105.4 33k 24.32 0.858 0.144 101.4 47k 23.67 0.791 0.225 255.2 168k
TABLE I: Average Metrics Across Datasets. Dataset-level averages on LLFF, NeRF Synthetic, and our MobileScan mobile-native dataset. Higher is better for PSNR/SSIM, and lower is better for LPIPS/Time. Best results are in bold.
Figure 4: Qualitative comparison on the LLFF dataset. Our method produces much sharper textures and more accurate thin structures (e.g., the leaves in Fern and the petals in Flower), closely matching the ground truth (GT).
Figure 5: Qualitative comparison on the NeRF Synthetic dataset. In these object-centric synthetic scenes, our method produces high-fidelity results.
Figure 6: Qualitative results on a mobile device across diverse datasets. We showcase real-time rendering screenshots from an iPhone 15, covering our self-collected scenes (rows 1-2), LLFF (row 3), and NeRF Synthetic (row 4). PocketGS consistently achieves high-fidelity reconstruction and sharp details across varying scene scales and capture conditions. The red labels indicate optimized point counts, demonstrating our method’s representation efficiency and robust generalization for practical on-device deployment.
IV-A Experimental Setup
IV-A1 Implementation Details and Experimental Settings
We implement PocketGS as a fully on-device system and evaluate it on an iPhone 15 (Apple A16). The entire process was written in Swift and accelerated by using the Apple Metal API to take advantage of the mobile GPU execution stack. Unless otherwise stated, all methods are evaluated under identical image resolution, color space, and metric implementations.
To reflect realistic mobile constraints, we use a fixed training budget for all methods. In particular, we optimize each model for a constant number of iterations (500 iterations) at a fixed view rendering resolution.
IV-A2 Baselines
[…]
3DGS-SFM-WK
We adopt the standard COLMAP-based Structure-from-Motion (SfM) pipeline on a workstation to estimate camera extrinsic/intrinsic parameters and reconstruct a sparse point cloud. This sparse geometric prior is then used to initialize and train a vanilla 3D Gaussian Splatting (3DGS) model, adhering to the original 3DGS training protocol.
3DGS-MVS-WK
For this workstation-based baseline, we leverage the standard COLMAP Multi-View Stereo (MVS) workflow to generate a dense point cloud as the geometric prior. Subsequently, a vanilla 3DGS model is initialized with this dense point cloud and trained following the default 3DGS optimization workflow.
We choose these two baselines because they (1) represent the standard reconstruction-first 3DGS pipeline, (2) span sparse SfM versus dense but costly MVS priors to separate prior quality from reconstruction overhead under matched budgets, and (3) provide a workstation reference that reflects PocketGS’s goal: competitive quality under strict on-device constraints without heavy offline reconstruction.
IV-A3 Datasets and Metrics
Datasets.
We evaluate on three representative datasets spanning synthetic and real captures: NeRF-Synthetic [24], LLFF [23], and MobileScan. NeRF-Synthetic provides object-centric synthetic scenes with accurate poses; LLFF consists of real forward-facing captures exhibiting real-world degradations; and MobileScan is our Phone-captured dataset designed to reflect mobile imaging characteristics, including motion blur, defocus, and exposure variation. For MobileScan, we use ARKit-provided intrinsics, poses, and sparse structure to emulate practical on-device inputs, enabling evaluation under noisy geometric priors and limited compute budgets. More details are in the supplement.
Metrics.
We report standard novel-view synthesis metrics, including PSNR (dB, $\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$). We additionally report end-to-end runtime to capture practical deployability. Specifically, we define the total runtime as
$T_{\text{total}}=T_{\text{geom}}+T_{\text{train}},$ (11)
where $T_{\text{geom}}$ includes geometry and pose acquisition as well as associated optimization, and $T_{\text{train}}$ denotes the 3DGS optimization time. For fair comparison, all methods are evaluated under the same training budget and at the same rendering resolution. Finally, we report Count (the number of Gaussians) to characterize the trade-off between representation capacity and computational efficiency.
IV-B Main Results: Quality and Efficiency Comparison
Table I summarizes averages under a strict, matched optimization budget. Across all datasets, PocketGS consistently delivers a strong quality–efficiency trade-off while maintaining compact or comparable Gaussian counts, indicating better representation efficiency.
LLFF Dataset.
LLFF stresses forward-facing captures and view extrapolation, where imperfect geometry can destabilize optimization. PocketGS achieves the best overall quality while remaining fast. Notably, 3DGS-MVS-WK performs poorly despite a larger Count, suggesting that a generic dense prior does not ensure good convergence in this regime. In contrast, PocketGS attains higher quality with fewer Gaussians, indicating better representation efficiency and conditioning under the same iteration budget.
Fig. 4 shows PocketGS better preserves fine, high-frequency geometry structures (e.g., leaves in Fern, petals in Flower), with fewer floating artifacts and cleaner boundaries. This is consistent with our refined geometry prior ($\mathcal{G}$) and robust anisotropic initialization ($\mathcal{I}$), which improve pose/structure consistency and reduce Gaussian drift during optimization.
NeRF-Synthetic Dataset.
On the geometry-clean NeRF Synthetic dataset, 3DGS-MVS-WK achieves the best metrics, as its dense prior matches the ground-truth geometry. PocketGS remains competitive (PSNR 24.32 vs. 24.47) while being $\sim 5.2\times$ faster in end-to-end runtime $T_{\text{total}}$ (101.4s vs. 532.1s), mainly by avoiding the costly offline dense reconstruction in 3DGS-MVS-WK. With a similar Count (47k vs. 50k), this speedup reflects a more streamlined pipeline and hardware-aligned splatting rather than reduced capacity.
Fig. 5 shows PocketGS preserves high-fidelity appearance and clean edges under the same budget, closely matching 3DGS-MVS-WK, demonstrating that PocketGS can substantially reduce capture-to-reconstruction latency without sacrificing quality when geometry is well-conditioned.
MobileScan Dataset.
On the challenging MobileScan dataset, PocketGS consistently outperforms 3DGS-SFM-WK, indicating that our mobile-native prior construction mitigates instability from sparse/noisy mobile inputs. Compared to 3DGS-MVS-WK, PocketGS achieves better perceptual quality (LPIPS 0.225 vs. 0.281) with much lower end-to-end time (255.2s vs. 534.5s), suggesting that heavyweight offline dense reconstruction is inefficient for mobile deployment. Notably, although PocketGS and 3DGS-MVS-WK converge to similar primitives, PocketGS renders more faithfully, implying improved placement rather than merely the quantity of Gaussians.
Fig. 3 shows that 3DGS-SFM-WK often yields blurry textures and structural collapse, consistent with its low Count (23k) and unreliable geometry. While 3DGS-MVS-WK starts from a dense prior, it still exhibits artifacts when Gaussians are overly concentrated; under a limited optimization budget, it cannot sufficiently re-balance spatial coverage, leading to ghosting and local inconsistency. PocketGS avoids this by surface-aligned, spatially balanced anisotropic initialization, producing sharper textures and higher structural fidelity closer to GT.
On-device Rendering and Deployment Perspective.
Beyond metrics, Fig. 6 demonstrates real-time rendering screenshots produced on an iPhone 15 across MobileScan, LLFF, and NeRF Synthetic scenes. The consistently sharp details across diverse scales and capture conditions indicate robust generalization in practical deployment. PocketGS maintains high perceptual quality under mobile memory and bandwidth constraints by avoiding unnecessary densification while preserving critical structures.
IV-C Memory Footprint
We profile peak memory usage on MobileScan to quantify the practicality of our on-device pipeline. Following our end-to-end definition, we report memory for two stages: (i) geometry prior construction and (ii) full 3DGS training. In all measurements, peak memory is recorded as the maximum resident memory observed during execution, and the statistics are aggregated over all MobileScan scenes.
Full 3DGS training reaches an average peak of 2.21 GB (range: 1.82–2.65 GB), reflecting the combined cost of storing the Gaussian parameters, intermediate buffers for rasterization, and optimizer states. Geometry prior construction peaks at 1.53 GB on average (range: 1.19–2.22 GB), dominated by temporary buffers used for feature processing, correspondence aggregation, and incremental updates of the prior representation. Importantly, across all scenes the peak memory remains below 3 GB, which fits within the on-device budget and leaves practical headroom for commodity smartphones to accommodate system overhead and concurrent background processes. We provide per-scene peak memory breakdowns for both stages in the supplemental material.
IV-D Ablation Study
We conduct an ablation study on the MobileScan dataset to quantify the contribution of each core component of PocketGS. The study follows the same experiment setup, and we summarize dataset averages in Table II.
[…]
PocketGS (Full) 23.67 0.791 0.225 255.2
w/o Initialization ($\mathcal{I}$) 22.49 0.7696 0.253 319.5
w/o Global BA 23.45 0.7517 0.232 251.1
w/o MVS 21.07 0.6461 0.414 124.8
TABLE II: Ablation Summary on MobileScan Dataset (Averages).
Figure 7: Ablation of information-gated frame subsampling. Our method can obtain better details information.
Information-Gated Frame Subsampling. Removing information-gated subsampling admits motion/defocus-degraded frames, corrupting the geometry prior and cascading into poorer initialization and blurrier reconstructions; the full pipeline better preserves sharp details (fig. 7).
GPU-native Global BA. Disabling Global BA yields a clear SSIM drop (from 0.791 to 0.7517) with minor PSNR/LPIPS changes, indicating reduced global coherence and pose consistency, while incurring negligible runtime overhead (251.1s vs. 255.2s).
Single-Reference Cost-Volume MVS. Removing lightweight MVS causes the largest quality degradation (PSNR 21.07, SSIM 0.6461, LPIPS 0.4137), showing dense geometric support largely sets the quality ceiling; although runtime drops to 124.8s, quality falls to the level of 3DGS-SFM-WK.
[…]
V Conclusion
We introduce PocketGS, a fully on-device 3DGS training paradigm that co-designs geometry priors ($\mathcal{G}$), prior-conditioned initialization ($\mathcal{I}$), and hardware-aligned splatting optimization ($\mathcal{T}$) to make high-fidelity 3D reconstruction practical under mobile compute and memory constraints. PocketGS achieves perceptual quality competitive with or superior to workstation baselines while meeting smartphone budgets (e.g., $\sim$4-minute end-to-end training and <3<3 GB peak memory on iPhone 15), enabling practical professional-grade mobile 3D content creation for mixed reality and digital twins.
Appendix A Implementation Details
In this section, we provide additional details regarding the three core operators of PocketGS: Geometry Prior Construction ($\mathcal{G}$), Prior-Conditioned Parameterization ($\mathcal{I}$), and Hardware-Aligned Splatting ($\mathcal{T}$).
A-A Geometry Prior Construction ($\mathcal{G}$)
Information-Gated Frame Selection
To ensure high-quality input under a limited time budget, we implement a real-time frame selection mechanism during the capture phase. We utilize a displacement gate $\tau=0.05$m to filter redundant frames. For frames passing the displacement gate, we employ a candidate window mechanism (maximum 8 frames or 0.25s) to select the sharpest frame. Sharpness is computed by calculating the gradient energy on a sparse $160\times 160$ grid of the luma plane, prioritizing frames with higher high-frequency content to mitigate motion blur.
GPU-Native Global Bundle Adjustment
Our global BA optimizes both camera poses and 3D point coordinates. We establish geometric constraints by extracting ORB/FAST features with BRIEF descriptors, accelerated via Metal. Feature matching is performed using Hamming distance with ratio tests and cross-checks. We implement an iterative refinement process (typically 3 rounds) that includes re-triangulation and outlier filtering based on reprojection error and triangulation angles. The optimization uses a Levenberg-Marquardt (LM) solver with Schur complement, fully resident on the GPU. To fix the gauge, we anchor the first frame’s 6DoF and the scale of the baseline between the first two frames.
Lightweight Single-Reference MVS
For dense reconstruction, we estimate per-frame depth maps using a plane-sweep algorithm. To optimize memory, we select the top-3 reference frames based on baseline distance and viewing angle but only utilize the best reference frame for cost volume construction. The depth search range is dynamically estimated using the 5%–95% quantiles of sparse point depths. We use Census transform for robust matching and Semi-Global Matching (SGM) for cost aggregation. The final dense point cloud is generated by back-projecting pixels with confidence $\geq 0.4$ and depth within 0.050.05m to 5.05.0m.
A-B Prior-Conditioned Parameterization ($\mathcal{I}$)
We initialize Gaussian primitives by leveraging the local surface statistics derived from the MVS point cloud. For each point, we estimate the local surface normal using PCA on its $k=16$ nearest neighbors. The scale parameters are initialized based on the average distance to the $k=3$ nearest neighbors. To seed anisotropic Gaussians, we set the scale along the normal direction to $0.3\times$ the tangential scales, effectively creating disc-like primitives that better align with the scene geometry.
A-C Hardware-Aligned Splatting ($\mathcal{T}$)
Our differentiable renderer is tailored for Tile-Based Deferred Rendering (TBDR) GPUs. We implement a multi-stage pipeline where alpha compositing is manually unrolled within the fragment shader. This allows us to cache intermediate states in tile memory, avoiding expensive global memory traffic. For backpropagation, we use index-mapped gradient scattering to ensure that gradients are correctly accumulated into the canonical parameter buffer despite the depth-sorted rendering order.
Appendix B PocketGS Mobile Application
To demonstrate the feasibility and practicality of our end-to-end on-device training system, we developed a dedicated mobile application, PocketGS, which integrates the entire pipeline from data capture to 3DGS training. The application is built on the Metal framework for high-performance GPU computation.
(a) Main Menu
(b) Splat Capture
(c) Process Capture
(d) Splat Training
Figure 8: Screenshots of the PocketGS mobile application, illustrating the end-to-end workflow.
The application workflow, as illustrated in Figure 8, is designed to be intuitive and efficient:
  • Main Menu (Fig. 8(a)): Provides access to the core functionalities: Splat Capture (data acquisition), View Point Cloud and View 3DGS (visualization), Splat Training (optimization), Import Dataset, and Process Capture (geometry prior construction).
  • Splat Capture (Fig. 8(b)): A guided capture interface that uses the information-gated frame selection mechanism to ensure high-quality input images (e.g., 50 frames) are collected in real-time, minimizing motion blur and redundancy.
  • Process Capture (Fig. 8(c)): This module manages captured scenes and executes the Geometry Prior Construction ($\mathcal{G}$) operator, including the GPU-native BA and lightweight MVS, to generate the initial PLY point cloud. Users can manage captured data and trigger the processing step.
  • Splat Training (Fig. 8(d)): This interface allows users to configure and initiate the on-device 3DGS training. Key features include selecting the input PLY file, choosing between Original Pose (ARKit pose) and Optimized Pose (BA-refined pose), and setting core hyperparameters (e.g., learning rates, iterations). The training process, which implements the $\mathcal{I}$ and $\mathcal{T}$ operators, runs entirely on the mobile GPU, with progress displayed in real-time.
The PocketGS app serves as a concrete proof-of-concept, demonstrating that high-quality 3DGS training can be achieved end-to-end on a commodity mobile device within minutes.
Figure 9: The scenes on MobileScan Dataset.
Appendix C MobileScan Dataset Details
The MobileScan dataset consists of 16 diverse indoor and outdoor scenes captured using iPhone 15. The statistics of the collected scenes are summarized in Table III. Each scene includes 50 high-resolution images with corresponding ARKit poses and sparse point clouds as baselines, as shown on 9.
Category Scenes Environment
Furniture Sofa,Table, Desk, Shelf Indoor
Objects Robot, Audio, Printer, Bag, Elxbox, audio. Indoor
Outdoor Pier, Shrub, Bench, Bike, Transformer, Chair, Flowerpot Outdoor
TABLE III: Statistics of the MobileScan Dataset.
The dataset covers various challenging scenarios, including:
  • Textureless Surfaces: Large uniform areas like the ”Sofa” and ”Desk” scenes.
  • Complex Geometry: Intricate structures in the ”Transformer” and ”Shrub” scenes.
  • Varying Lighting: Outdoor scenes like ”Pier” with natural lighting transitions.
Appendix D Additional Quantitative Results
We provide per-scene metrics for our MobileScan dataset, NeRF Synthetic dataset, and LLFF dataset. These tables serve as the full, unaggregated data supporting the main paper’s claims on performance metrics (PSNR, SSIM, LPIPS) and efficiency (Time, Count).
Scene 3DGS-SFM-WK 3DGS-MVS-WK PocketGS (Ours)
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s) $\downarrow$ Count PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s) $\downarrow$ Count PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s) $\downarrow$ Count
Pier 20.23 0.498 0.495 156.0 35800 20.54 0.695 0.312 519.6 200936 20.29 0.625 0.329 212.8 176804
Chair 22.13 0.542 0.421 174.0 39420 21.34 0.699 0.270 555.1 180664 22.39 0.670 0.256 240.7 176970
Flowerpot 20.93 0.595 0.424 168.0 32415 21.00 0.730 0.281 486.9 242616 21.97 0.725 0.225 265.7 241976
Shrub 18.88 0.507 0.481 204.0 37476 18.53 0.582 0.427 546.7 98616 18.94 0.609 0.331 230.0 174278
Transformer 20.49 0.542 0.465 162.0 27678 19.66 0.636 0.360 546.7 122117 21.05 0.636 0.315 209.7 126222
Sofa 14.26 0.611 0.558 28.6 38 25.32 0.870 0.222 525.4 312949 24.90 0.873 0.192 389.4 289848
Table 19.71 0.754 0.382 78.0 18873 15.01 0.727 0.365 401.9 113738 22.85 0.852 0.194 184.6 179940
Robot 22.90 0.807 0.285 114.0 20795 22.21 0.862 0.207 581.7 209226 24.64 0.850 0.167 291.9 194881
Desk 20.31 0.754 0.313 72.0 20735 17.42 0.750 0.279 491.9 111202 22.03 0.838 0.172 192.8 108018
Elebox 26.17 0.922 0.197 22.4 4410 21.11 0.899 0.181 575.8 52654 27.75 0.936 0.136 255.1 37182
audio 26.05 0.815 0.330 150.0 28628 23.97 0.852 0.263 520.0 147733 27.69 0.857 0.166 270.2 155278
bag 23.71 0.855 0.313 108.0 22584 21.35 0.865 0.260 533.0 170320 24.79 0.859 0.210 320.4 179101
fireExtin 23.94 0.879 0.274 96.0 18162 22.09 0.886 0.239 592.0 146224 26.40 0.896 0.170 279.8 123325
printer 22.56 0.804 0.376 59.8 10368 20.08 0.834 0.277 542.8 203770 26.30 0.882 0.202 282.6 196896
shelf 20.39 0.761 0.418 24.8 4860 21.17 0.811 0.313 518.8 77940 25.40 0.872 0.226 224.4 75872
umbrella 24.29 0.813 0.341 47.4 9736 21.98 0.845 0.273 581.8 179388 23.30 0.826 0.235 261.9 173587
Average 21.16 0.687 0.398 112.8 23486 20.85 0.781 0.281 534.5 165367 23.67 0.791 0.225 255.2 168009
TABLE IV: MobileScan dataset comparison. Best results are highlighted in bold
Scene 3DGS-SFM-WK 3DGS-MVS-WK PocketGS (Ours)
[…]
Chair 22.11 0.869 0.154 90.0 14433 25.37 0.957 0.054 692.4 60859 23.70 0.870 0.089 85.7 49857
Drums 18.78 0.762 0.268 53.8 12842 21.27 0.861 0.154 570.1 60640 20.97 0.775 0.183 130.1 58511
Ficus 19.99 0.812 0.212 44.2 1511 20.36 0.804 0.233 144.3 5674 22.94 0.879 0.134 33.5 5638
Hotdog 23.13 0.829 0.236 84.0 9703 29.09 0.966 0.050 571.3 40686 28.07 0.940 0.094 78.8 38583
Lego 22.02 0.795 0.228 132.0 29883 25.47 0.930 0.092 770.3 73821 24.92 0.872 0.124 130.5 74572
Materials 21.21 0.779 0.254 120.0 6934 22.45 0.843 0.170 414.8 51456 22.85 0.842 0.158 117.4 52137
Mic 23.62 0.859 0.205 31.7 5879 27.11 0.936 0.066 519.5 39553 26.24 0.901 0.120 81.0 36082
Ship 23.14 0.696 0.386 114.0 15006 24.60 0.798 0.202 574.5 69482 24.87 0.781 0.248 153.9 64208
Average 21.75 0.800 0.243 83.7 12024 24.47 0.887 0.128 532.1 50271 24.32 0.858 0.144 101.4 47449
TABLE V: NeRF Synthetic Dataset comparison.
Scene 3DGS-SFM-WK 3DGS-MVS-WK PocketGS (Ours)
[…]
Fern 21.11 0.649 0.415 21.0 11436 21.36 0.707 0.335 118.2 38816 25.75 0.849 0.182 107.4 29220
Flower 22.97 0.701 0.328 114.0 22440 20.53 0.632 0.381 316.4 34914 25.06 0.818 0.187 92.0 28523
Fortress 26.51 0.744 0.321 180.0 29365 23.59 0.762 0.277 551.9 69762 27.15 0.901 0.129 91.6 67120
Horns 21.19 0.674 0.412 126.0 22335 17.42 0.626 0.432 530.9 38403 23.39 0.806 0.233 148.6 35996
Leaves 15.36 0.352 0.600 246.0 9358 15.28 0.327 0.588 190.4 20315 15.77 0.456 0.501 60.8 9358
Orchids 18.85 0.562 0.401 66.0 21968 18.11 0.593 0.353 141.1 66975 21.73 0.772 0.183 138.9 55320
Room 21.70 0.764 0.400 15.3 4737 21.34 0.787 0.358 203.4 27610 26.18 0.897 0.168 75.6 19273
Trex 20.42 0.684 0.360 96.0 21645 18.62 0.664 0.376 452.4 30104 23.33 0.830 0.194 128.6 22349
Average 21.01 0.641 0.405 108.0 17910 19.53 0.637 0.387 313.1 40862 23.54 0.791 0.222 105.4 33395
TABLE VI: LLFF dataset comparison.
D-A MobileScan Dataset Comparison
The detailed per-scene quantitative results on our d MobileScan dataset are presented in Table IV. PocketGS consistently outperforms 3DGS-SFM-WK and 3DGS-MVS-WK in terms of perceptual quality (LPIPS) while maintaining a significantly lower training time than the dense baseline.
Metric audio printer fireExtin flowerpot Elebox desk transformer shrub bench sofa Keybroad bike Pier bag robot umbrella chair table shelf Avg.
MVS Peak Mem. (GB) 1.55 1.50 1.55 1.48 1.38 1.19 1.48 1.64 2.22 1.39 1.67 1.50 1.73 1.42 1.57 1.56 1.49 1.24 1.47 1.53
Training Peak Mem. (GB) 2.65 2.62 2.50 2.49 2.44 2.41 2.29 2.27 2.20 2.15 2.15 2.12 2.03 2.02 1.98 1.95 1.94 1.91 1.82 2.21
TABLE VII: Per-scene peak memory.
Metric Elebox Pier audio bag bench bike chair desk fireExtin flowerpot printer robot shelf shrub sofa table transformer umbrella Avg.
PSNR (dB) 27.03 20.16 20.93 22.72 21.97 23.02 22.09 22.07 26.46 20.98 24.83 21.35 24.99 18.80 22.05 22.89 20.91 21.55 22.49
SSIM 0.9335 0.6048 0.8138 0.8303 0.6859 0.7185 0.6514 0.8385 0.8961 0.6836 0.8555 0.7934 0.8647 0.5889 0.8203 0.8522 0.6255 0.7955 0.7696
LPIPS 0.1400 0.3572 0.1859 0.2628 0.3207 0.2553 0.2760 0.1726 0.1781 0.2735 0.2390 0.2259 0.2481 0.3588 0.2706 0.1935 0.3316 0.2699 0.2533
Time (s) 369.1 325.6 363.0 335.8 350.4 327.7 306.8 186.7 429.2 386.9 352.7 283.0 318.3 331.5 329.3 172.6 297.7 284.5 319.5
TABLE VIII: Per-scene metrics w/o initialization (removing prior-conditioned anisotropic seeding in $\mathcal{I}$).
Metric Elebox Pier audio bag bench bike chair desk fireExtin flowerpot printer robot shelf shrub sofa table transformer umbrella Avg.
PSNR (dB) 27.61 20.21 26.95 24.47 22.40 23.03 22.30 22.02 26.45 21.67 26.10 24.38 24.85 18.82 24.49 22.52 20.80 23.01 23.45
SSIM 0.8975 0.5910 0.8156 0.8201 0.6869 0.6832 0.6318 0.8014 0.8548 0.6843 0.8401 0.8048 0.8282 0.5774 0.8283 0.8107 0.5915 0.7825 0.7517
LPIPS 0.1350 0.3253 0.1743 0.2172 0.2857 0.2595 0.2639 0.1758 0.1821 0.2326 0.2093 0.1869 0.2367 0.3311 0.1928 0.2026 0.3285 0.2437 0.2324
Time (s) 230.7 212.2 280.5 311.5 259.9 231.9 233.5 181.3 285.0 254.2 282.5 263.2 232.0 230.6 352.7 176.7 238.5 263.0 251.1
TABLE IX: Per-scene metrics w/o global BA (removing GPU-native global BA in $\mathcal{G}$).
Metric Elebox Pier audio bag bench bike chair desk fireExtin flowerpot printer robot shelf shrub sofa table transformer umbrella Avg.
PSNR (dB) 23.10 18.25 24.43 23.67 19.84 21.03 19.46 20.12 23.35 19.56 23.38 22.66 21.09 16.98 21.96 19.61 18.91 21.92 21.07
SSIM 0.9029 0.3181 0.7425 0.8163 0.4339 0.5333 0.3467 0.7574 0.8413 0.5089 0.8195 0.7925 0.7845 0.2820 0.7690 0.7533 0.4489 0.7795 0.6461
LPIPS 0.2113 0.6392 0.3905 0.3201 0.5653 0.4897 0.5609 0.2855 0.3064 0.4894 0.3259 0.2600 0.3583 0.6229 0.3898 0.3478 0.5480 0.3366 0.4137
Time (s) 122.2 93.8 186.7 142.1 101.4 110.7 129.0 95.9 178.5 118.9 149.3 112.2 142.3 88.2 136.2 89.9 111.0 137.5 124.8
TABLE X: Per-scene metrics w/o MVS (removing lightweight single-reference MVS in $\mathcal{G}$).
D-B NeRF Synthetic and LLFF Datasets
We also evaluate PocketGS on standard benchmarks to demonstrate its generalization capability. The per-scene results for the NeRF Synthetic dataset are shown in Table V, and the LLFF dataset results are in Table VI.
Appendix E Memory Analysis
We analyze the peak memory usage of PocketGS during the MVS and training phases. The detailed per-scene memory usage is provided in Table VII, demonstrating that our system operates well within the memory budget of commodity mobile devices.
Appendix F Ablation Study
We conduct ablation studies to evaluate the impact of our core components. The following tables provide the detailed per-scene metrics for each ablation configuration, serving as the full data to support the aggregated results presented in the main paper.
F-A Ablation: w/o Initialization ($\mathcal{I}$)
This configuration removes the prior-conditioned anisotropic seeding in $\mathcal{I}$, reverting to isotropic initialization. The detailed per-scene results in Table VIII show a significant drop in performance, highlighting the importance of geometry-aware initialization for fast convergence.
F-B Ablation: w/o Global BA
This configuration removes the GPU-native global Bundle Adjustment (BA) in $\mathcal{G}$, relying solely on the initial ARKit poses. The detailed per-scene results in Table IX indicate that the global BA is crucial for refining the coarse mobile poses, leading to better reconstruction quality.
F-C Ablation: w/o MVS
This configuration removes the lightweight single-reference MVS in $\mathcal{G}$, relying only on the sparse point cloud from ARKit for initialization. The significant drop in all metrics, as shown in Table X, demonstrates the necessity of a dense, geometry-faithful prior for achieving high perceptual quality within a limited training budget.

### v3

@@
 Training 3DGS on device introduces substantial challenges, as it must deliver high perceptual quality under the limited computational resources and strict thermal constraints of mobile platforms.
-Consequently, the training process must meet on-device constraints: 1) completion within a few minutes ( $\textless 5$ min), 2) respecting the hardware-available memory budget ( $\textless 3$ GB).
+Consequently, the training process must meet on-device constraints: 1) completion within a few minutes (e.g., <5<5 min), 2) respecting the hardware-available memory budget (e.g., <3<3 GB).
 These constraints require rethinking 3DGS under resource-limited regimes beyond conventional resource-unconstrained assumptions.
@@
 Existing designs assume ample time and memory across all stages, but on-device settings break these assumptions, yielding three tightly coupled contradictions.
+21
 Input-Recovery Contradiction.
@@
 Thus, on-device 3DGS requires geometry-conditioned initialization to reduce optimization burden, thereby facilitating stable convergence.
-Hardware-Differentiability Contradiction.
-On-device 3DGS is at odds with mobile execution: differentiable rendering requires storing intermediates for backpropagation, but mobile pipelines restrict gradient access through opaque blending, suffer from costly GPU–CPU synchronization in optimization, and incur heavy data movement and state misalignment from depth-ordered alpha blending.
-Therefore, on-device 3DGS requires a mobile-oriented differentiable rendering framework that bounds and replays only the necessary intermediates for Gaussian optimization.
+Hardware–Differentiability Mismatch.
+On-device 3DGS conflicts with mobile execution: while differentiable splatting needs per-step alpha-compositing states for backpropagation, mobile fixed-function pipelines expose only the final blended color and conceal the depth-ordered accumulation needed to route gradients.
+Recovering these states via framebuffer readbacks, full buffer clears, or backward-time re-enumeration of per-pixel splat sequences is bandwidth-prohibitive and poorly aligned with mobile GPU state.
+Thus, on-device 3DGS calls for a mobile-oriented differentiable rendering framework that explicitly unrolls compositing and replays only the minimal cached intermediates required for Gaussian optimization.
 In this work, we present PocketGS, a 3DGS training paradigm that achieves high-perceptual modeling under on-device constraints via three co-designed operators (Fig.
-LABEL:top,2).
+1,2).
 To address the Input-Recovery Contradiction, the geometry-prior construction operator $\mathcal{G}$ builds a geometry-faithful prior, providing the required geometric guidance without costly in-loop Gaussian densification beyond the on-device envelope.
 To resolve the Initialization-Convergence Contradiction, the prior-conditioned Gaussian parameterization operator $\mathcal{I}$ uses this prior to shape initial Gaussians, improving conditioning for stable, iteration-limited optimization and reducing degradation from noise-sensitive initialization.
-Finally, to overcome the Hardware-Differentiability Contradiction, the hardware-aligned splatting optimization operator $\mathcal{T}$ makes differentiable rendering practical on mobile GPUs, preserving correct gradients and stable updates without prohibitive memory traffic or CPU–GPU synchronization.
+Finally, to overcome the Hardware-Differentiability Contradiction, the hardware-aligned splatting optimization operator $\mathcal{T}$ makes differentiable rendering practical on mobile GPUs, preserving correct gradients and stable updates without prohibitive memory traffic.
 Extensive experiments show that PocketGS achieves perceptual quality superior to mainstream workstation 3DGS baselines, while operating within on-device constraints.
@@
   • We develop a hardware-aligned splatting optimization operator $\mathcal{T}$ that renders differentiable splatting practical on mobile GPUs, supporting accurate gradients and stable updates within hardware memory and bandwidth constraints.
-  • We validate that, while satisfying on-device constraints, PocketGS achieves competitive perceptual quality relative to standard workstation 3DGS pipelines on publicly available benchmarks and our self-collected MobileScan datasets.
+We empirically show that PocketGS satisfies strict on-device runtime and peak-memory budgets while achieving competitive perceptual quality compared with standard workstation 3DGS pipelines, on both publicly available benchmarks and our self-collected MobileScan dataset.
 Figure 2: Overview of the PocketGS framework.
@@
 ($\mathcal{I}$) Prior-Conditioned Parameterization seeds anisotropic Gaussians by estimating local surface statistics (normals $\mathbf{s}_{n}$ and scales) to front-load structure discovery via disc-like covariance seeding.
-($\mathcal{T}$) Hardware-Aligned Splatting implements a mobile-native differentiable renderer using unrolled alpha-compositing ($S={C_{in},\alpha}$) and index-mapped gradient scattering to ensure stable backpropagation within tight mobile memory bounds of the canonical parameter buffer ($\mu,\Sigma,c,\alpha$).
+($\mathcal{T}$) Hardware-Aligned Splatting implements a mobile-native differentiable renderer using unrolled alpha-compositing ($S={id,C_{in},\alpha}$) and index-mapped gradient scattering to ensure stable backpropagation within tight mobile memory bounds of the canonical parameter buffer ($\mu,\Sigma,c,\alpha$).
 II Related Work
@@
 III The PocketGS Paradigm
-PocketGS enables on-device 3DGS training: given a short mobile capture $I_{t}$}, it optimizes a Gaussian scene $\Theta$ under strict runtime and peak-memory budgets on tile-based mobile GPUs.
-It co-designs three coupled operators: $\mathcal{G}$ derives a compact geometry prior $\mathcal{P}$ from $I_{t}$}, $\mathcal{I}$ uses $\mathcal{P}$ to initialize well-conditioned Gaussians $\Theta_{0}$, and $\mathcal{T}$ performs hardware-aligned differentiable splatting to keep training GPU-resident with correct gradients.
-Formally, $\mathcal{G}:(\{I_{t}\},\{\hat{T}_{t}\})\mapsto(\{T_{t}\},\mathcal{P}),\ \mathcal{I}:\mathcal{P}\mapsto\Theta_{0},$ and $\mathcal{T}:(\Theta_{0},\{I_{t},T_{t}\})\mapsto\Theta^{*}$, balancing perceptual quality with on-device constraints.
+PocketGS enables on-device 3D Gaussian Splatting (3DGS) training under strict runtime and peak-memory constraints.
+Given a sequence of video frames $\{I_{t}\}_{t=1}^{n}$ and their corresponding coarse camera poses $\{\hat{T}_{t}\}_{t=1}^{n}$ obtained from mobile tracking systems (ARKit or ARCore), PocketGS comprises three tightly coupled operators.
+Specifically, the geometry operator $\mathcal{G}$ refines the coarse poses $\{\hat{T}_{t}\}$ into accurate ones $\{{T}_{t}\}$, and extracts a dense point cloud $P$, which serves as a reliable geometric prior.
+The initialization operator $\mathcal{I}$ maps $P$ to a well-conditioned 3D Gaussian initialization $\Theta_{0}$.
+Finally, the training operator $\mathcal{T}$ performs hardware-aligned differentiable rendering to optimize $\Theta_{0}$ into the final Gaussian representation $\Theta^{*}$.
+Formally, these operators are defined as: $\mathcal{G}:(\{I_{t},\hat{T}_{t}\})\mapsto(\{T_{t}\},P)$, $\mathcal{I}:P\mapsto\Theta_{0}$, and $\mathcal{T}:(\Theta_{0},\{I_{t},T_{t}\})\mapsto\Theta^{*}$.
 III-A Geometry Prior Construction under Resource Budgets
-To resolve the Input-Recovery Contradiction, Operator $\mathcal{G}$ constructs a low-noise and low-redundant 3DGS-aligned prior within mobile compute and memory constraints.
-Treating BA (Bundle Adjustment) and MVS (Multi-View Stereo) as core prior-construction primitives, their outputs directly condition the 3DGS optimization, mitigating over-reliance on excessive in-loop Gaussian densification.
+Prior 3DGS approaches rely on prolonged optimization with training-time densification to recover scenes from noisy and sparse point inputs.
+On mobile devices, this paradigm often yields unstable geometry recovery and substantial computational overhead, making it impractical under on-device constraints.
+We instead leverage the mobile capture $\{I_{t},\hat{T}_{t}\}$ to build a dense, low-noise, and low-redundancy geometry prior by applying a geometry operator $\mathcal{G}$, which maps $(\{I_{t},\hat{T}_{t}\})$ to refined poses $\{T_{t}\}$ and a compact geometric representation $P$.
+Conditioning the subsequent optimization on $(\{T_{t}\},P)$ removes the need for training-time densification, resulting in more stable convergence and improved computational efficiency.
 III-A1 Information-Gated Frame Subsampling
-To reduce input noise and provide high-quality photometric ground truth for 3DGS training, PocketGS adopts a real-time three-stage gating mechanism that admits only frames with incremental geometric utility, maximizing information yield per compute unit.
+Video frames captured during mobile scanning are often affected by motion blur, which degrades geometric reconstruction and undermines the photometric optimization of 3DGS.
+To mitigate this issue, we subsample the input sequence using an information-gated keyframe selection strategy based on viewpoint change and distance variation, retaining only sharp and high-utility frames while filtering out blurred and redundant ones.
 Displacement Gate.
-We ensure sufficient parallax for robust triangulation by computing the camera displacement $d$ between the current frame and the last selected keyframe, $d=\|\mathbf{t}_{curr}-\mathbf{t}_{last}\|_{2}$.
+We ensure sufficient parallax for robust triangulation by computing the camera displacement $d$ between the current frame $\mathbf{t}_{curr}$ and the last selected keyframe $\mathbf{t}_{last}$, $d=\|\mathbf{t}_{curr}-\mathbf{t}_{last}\|_{2}$.
 A frame is considered only if $d\geq\tau_{d}$, where $\tau_{d}=0.05m$, suppressing redundant viewpoints while maintaining tracking stability.
@@
 We estimate the sharpness score $S$ by summing absolute luma differences over a sparsely sampled grid $\Omega$:
-$S=\frac{1}{|\Omega|}\sum_{(x,y)\in\Omega}\left(|I(x+\Delta,y)-I(x,y)|+|I(x,y+\Delta)-I(x,y)|\right)$ (1)
+$\small S=\frac{1}{|\Omega|}\sum_{(x,y)\in\Omega}\left(|I(x+\Delta,y)-I(x,y)|+|I(x,y+\Delta)-I(x,y)|\right)$ (1)
 where $\Delta$ is the step size. This score correlates with high-frequency content and favors sharp frames.
 Candidate Windowing.
-We select the locally optimal frame from a short temporal window (8 frames or 250ms).
+We select the locally optimal frame from a short temporal window (8 frames).
 A candidate replaces the current best only if $S_{new}>(1+r)S_{best}$, with $r=0.05$.
 This commits the sharpest representative per viewpoint and bounds downstream optimization workload.
-Keyframe admission is explicitly bounded by the three gates, reducing downstream BA/MVS workload roughly proportional to admitted keyframes rather than raw capture length, thereby constraining both runtime and peak intermediate storage.
 III-A2 GPU-Native Global BA as Mobile MAP Refinement
-To address mobile input imprecision, PocketGS implements GPU-native global BA to jointly optimize poses and 3D points, delivering 3DGS-required precision without over-densification.
+Initial camera poses provided by mobile tracking systems (e.g., ARKit and ARCore) are often noisy, which degrades both geometric reconstruction and the photometric optimization of 3DGS.
+To address this issue, we perform GPU-native global BA on-device to refine poses, thereby providing accurate pose estimates for subsequent dense point recovery and 3DGS optimization.
 The robustified reprojection objective for PocketGS:
-$\min_{\{\mathbf{T}_{i}\},\{\mathbf{P}_{j}\}}\sum_{i,j}\rho\left(\|\pi(\mathbf{T}_{i},\mathbf{P}_{j})-\mathbf{p}_{ij}\|^{2}_{\Sigma_{ij}}\right)$ (2)
-where $\pi$ is the projection function, $\mathbf{T}_{i}\in SE(3)$ is the pose of camera $i$, $\mathbf{P}_{j}\in\mathbb{R}^{3}$ is the 3D position of point $j$, $\mathbf{p}_{ij}\in\mathbb{R}^{2}$ is the observed 2D projection, and $\rho$ is the robust Huber loss.
+$\min_{\{\mathbf{\hat{T}}_{i}\},\{\mathbf{P}_{j}\}}\sum_{i,j}\rho\left(\|\pi(\mathbf{\hat{T}}_{i},\mathbf{P}_{j})-\mathbf{p}_{ij}\|^{2}_{\Sigma_{ij}}\right)$ (2)
+where $\pi$ is the projection function, $\mathbf{\hat{T}}_{i}\in SE(3)$ is the pose of camera $i$, $\mathbf{P}_{j}\in\mathbb{R}^{3}$ is the 3D position of point $j$, $\mathbf{p}_{ij}\in\mathbb{R}^{2}$ is the observed 2D projection, and $\rho$ is the robust Huber loss.
 Scale-Aware Gauge Fixing.
@@
 We exploit the sparse block structure via a full GPU Schur complement.
-Partitioning parameters into camera pose (c) and point (p) blocks yields the reduced camera system:
-$(\mathbf{H}_{cc}-\mathbf{H}_{cp}\mathbf{H}_{pp}^{-1}\mathbf{H}_{pc})\Delta_{c}=\mathbf{b}_{c}-\mathbf{H}_{cp}\mathbf{H}_{pp}^{-1}\mathbf{b}_{p}$ (3)
+Partitioning parameters into camera pose (t) and point (p) blocks yields the reduced camera system:
+$(\mathbf{H}_{tt}-\mathbf{H}_{tp}\mathbf{H}_{pp}^{-1}\mathbf{H}_{pt})\Delta_{c}=\mathbf{b}_{t}-\mathbf{H}_{tp}\mathbf{H}_{pp}^{-1}\mathbf{b}_{p}$ (3)
 Since $\mathbf{H}_{pp}$ is block-diagonal with independent $3\times 3$ blocks per point, $\mathbf{H}_{pp}^{-1}$ is computed in parallel across points on the GPU.
-After solving for $\Delta_{c}$, we recover $\Delta_{p}$ by back-substitution.
+After solving for $\Delta_{t}$, we recover $\Delta_{p}$ by back-substitution.
 Iterative Geometric Refinement.
@@
 This purification reduces outliers and stabilizes dense reconstruction.
-Schur complement reduces solving from the full joint system to a smaller camera system, while $\mathbf{H}_{pp}^{-1}$ is computed in parallel; the design keeps memory dominated by sparse observations rather than dense matrices, making global refinement feasible on-device.
 III-A3 Single-Reference Cost-Volume MVS
-To address mobile input sparsity, PocketGS’s lightweight MVS generates a dense geometry-faithful point cloud, providing structural richness for Gaussian initialization without over-densification.
+Point clouds provided by mobile tracking systems are typically sparse, which necessitates costly densification during 3DGS training.
+To address this limitation, our lightweight MVS reconstructs a dense and geometry-faithful point cloud for Gaussian initialization, thereby reducing over-densification under mobile budgets.
 Probabilistic Depth Range Estimation.
@@
 III-B Prior-Conditioned Gaussian Parameterization
-To ensure stable convergence under a strict mobile iteration budget, operator $\mathcal{I}$ improves early-stage conditioning by replacing isotropic initialization with prior-conditioned, surface-aligned anisotropic Gaussians, thereby injecting structural cues from $\mathcal{G}$ and reducing reliance on photometric signals to infer geometry.
+Prior methods initialize points as isotropic Gaussians based solely on inter-point distances, resulting in arbitrary parameter initialization that requires unnecessary updates and consequently slow convergence.
+In contrast, we exploit point-cloud surface geometry $P$ to estimate normals and initialize surface-aligned anisotropic Gaussians via an initialization operator $\mathcal{I}$, which maps $P$ to the initial Gaussian parameters $\Theta_{0}$.
+This structure-consistent starting state $\Theta_{0}=\mathcal{I}(P)$ improves conditioning and accelerates optimization under tight on-device budgets.
 III-B1 Local Surface Statistics Estimation
-To inject geometric prior knowledge into initialization, we estimate local surface statistics for each point $\mathbf{p}_{i}$ in the dense MVS point cloud via KNN (K=16).
-The neighborhood covariance matrix is computed as:
-$\mathbf{C}=\frac{1}{K}\sum_{k=1}^{K}(\mathbf{p}_{k}-\bar{\mathbf{p}})(\mathbf{p}_{k}-\bar{\mathbf{p}})^{T}$ (5)
-where $\bar{\mathbf{p}}_{i}$ is the centroid of the neighborhood.The eigenvector corresponding to $\mathbf{C}_{i}$’s smallest eigenvalue robustly estimates the local surface normal $\mathbf{n}_{i}$.
-This GPU-parallelized process provides critical structural cues for informed initialization.
-Fixed local neighborhood size ensures embarrassingly parallel computation with bounded per-point memory, aligning with PocketGS’s mobile constraints.
-III-B2 Disc-Like Covariance Seeding
-We initialize each Gaussian as a thin, disc-like ellipsoid tangent to the estimated surface, directly addressing the contradiction by replacing naive isotropic initialization with geometry-conditioned anisotropy.
-The tangential scale $s_{t}$ is derived from local point density, estimated as the average distance to the K=3 nearest neighbors:
-$s_{t}=\frac{1}{3}\sum_{k=1}^{3}\|\mathbf{p}-\mathbf{p}_{k}\|$ (6)
-We set the normal-direction scale $s_{n}$ as a fraction of tangential scale:
-$s_{n}=s_{t}\times r_{normal}$ (7)
-where $r_{normal}=0.3$.
-Scale parameters are optimized in log-space, $\mathbf{s}_{log}=\log(\mathbf{s})$, ensuring positivity and numerical stability.
-The initial rotation $\mathbf{q}$ aligns the Gaussian local z-axis with the estimated surface normal $\mathbf{n}_{i}$.
-Opacity is initialized to a fixed low value (logit of 0.1).
-This geometry-conditioned anisotropy supplies strong priors on surface orientation and extent, accelerating convergence under limited iterations.
+To extract geometric priors from the mobile capture, we estimate local surface statistics for each point $\mathbf{p}_{i}$ in the dense MVS point cloud using a fixed-size KNN neighborhood (K=16K{=}16).
+Let $\{\mathbf{p}_{i,k}\}_{k=1}^{K}$ denote the $K$ nearest neighbors of $\mathbf{p}_{i}$, with centroid
+$\bar{\mathbf{p}}_{i}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{p}_{i,k}.$ (5)
+We compute the local covariance as
+$\mathbf{C}_{i}=\frac{1}{K}\sum_{k=1}^{K}(\mathbf{p}_{i,k}-\bar{\mathbf{p}}_{i})(\mathbf{p}_{i,k}-\bar{\mathbf{p}}_{i})^{\mathsf{T}}.$ (6)
+The eigenvector of $\mathbf{C}_{i}$ associated with its smallest eigenvalue provides an estimate of the local surface normal $\mathbf{n}_{i}$.
+This per-point, fixed-$K$ procedure is embarrassingly parallel with bounded memory, making it well-suited for GPU execution under PocketGS’s mobile constraints.
 III-C Hardware-Aligned Differentiable Splatting Optimization
-Operator $\mathcal{T}$ rewrites the 3DGS training loop for GPU-resident optimization on mobile GPUs under tight memory/bandwidth budgets, reducing intermediate storage and synchronization overhead while enabling differentiable blending.
-By unrolling compositing and decoupling parameter/state layout from rendering order, it supports end-to-end on-device training with correct, stable gradients.
+On mobile GPUs, on-device 3D Gaussian Splatting (3DGS) training is bandwidth-bound: fixed-function blending reveals only the final color, hiding the intermediate compositing states needed for backpropagation, while auxiliary-buffer clears incur heavy memory traffic.
+We therefore align optimization with the mobile pipeline by unrolling front-to-back compositing into an explicit differentiable operator and caching a compact forward replay trace.
+This defines a GPU-resident map $\mathcal{T}:(\Theta_{0},\{I_{t},T_{t}\})\mapsto\Theta^{*}$ that delivers correct, stable gradients without framebuffer readbacks or backward-time reconstruction of per-pixel splat sequences.
 III-C1 Unrolled Alpha-Compositing with Forward Replay Cache
-To address hidden intermediate states, PocketGS unrolls alpha-compositing into an explicit differentiable pipeline with cached intermediates. Standard front-to-back compositing follows:
+Mobile fixed-function blending returns only the final composited color, while 3DGS backpropagation requires the intermediate accumulator state at each compositing step.
+Moreover, the set and order of Gaussians contributing to a pixel are not accessible from the rendering pipeline, making a backward pass that replays the original per-pixel splat list infeasible or prohibitively expensive.
+We address this by (1) unrolling the front-to-back compositing into an explicit computation graph and (2) caching a compact forward replay trace per pixel for backward traversal.
+Explicit Computation Graph.
+For depth-sorted visible Gaussians, we explicitly perform front-to-back compositing:
 $C_{out}=C_{in}(1-\alpha)+\alpha c,$ (7)
-where $C_{in}$ is the pre-composite color, and $(\alpha,c)$ are the opacity and color of the current Gaussian fragment.
-In a mobile rendering pipeline, hardware blending returns only $C_{out}$ while hiding intermediate factors, making correct gradient propagation unavailable without costly framebuffer reads.
-Explicit Computation Graph.
-We process visible Gaussians in depth-sorted batches. Each batch explicitly takes the previous accumulator as input, exposing the intermediates necessary for the backward pass:
-$\frac{\partial\mathcal{L}}{\partial C_{in}}=\frac{\partial\mathcal{L}}{\partial C_{out}}(1-\alpha),\quad\frac{\partial\mathcal{L}}{\partial\alpha}=\frac{\partial\mathcal{L}}{\partial C_{out}}(c-C_{in}).$ (9)
-Bandwidth-Efficient Cache.
-To minimize memory bandwidth pressure-a critical constraint for mobile energy efficiency-we introduce a forward replay cache.
-Instead of caching full per-pixel splat lists, we store a minimal set $S=\{C_{in},\alpha\}$ and use a per-pixel counter buffer.
-By resetting only the $O(WH)$ counter rather than the entire $O(WHK_{max})$ cache, we significantly reduce the bandwidth overhead during the backward traversal.
+where $C_{in}$ is the accumulator input and $(\alpha,c)$ are the opacity and color of the current Gaussian fragment. The backward pass follows the same unrolled chain:
+$\frac{\partial\mathcal{L}}{\partial C_{in}}=\frac{\partial\mathcal{L}}{\partial C_{out}}(1-\alpha),\qquad\frac{\partial\mathcal{L}}{\partial\alpha}=\left\langle\frac{\partial\mathcal{L}}{\partial C_{out}},\,(c-C_{in})\right\rangle,$ (8)
+where $\langle\cdot,\cdot\rangle$ denotes the per-channel inner product.
+Bandwidth-Efficient Forward Replay Cache.
+To avoid backward-time reconstruction of per-pixel splat sequences, we record only the intermediates necessary to replay gradients:
+$S=\{\mathrm{id},\,C_{in},\,\alpha\},$ (9)
+where $\mathrm{id}$ is the (sorted) splat identifier used to scatter gradients back to canonical parameter buffers.
+A per-pixel counter $\mathrm{count}(u)$ stores the number of valid cached entries for pixel $u$.
+Each iteration resets only the $O(WH)$ counters, rather than clearing the full $O(WHK_{\max})$ cache.
+During backpropagation, we traverse only
+$k\in[0,\mathrm{count}(u)-1],$
+treating stale cache entries beyond $\mathrm{count}(u)$ as invalid, thereby eliminating bandwidth-heavy full-cache clears while preserving gradient correctness.
 III-C2 Index-Mapped Gradient Scattering
-Maintaining optimizer consistency during depth-sorted rendering typically requires physical memory reordering, which incurs heavy data movement.
-We resolve this by decoupling the sorted view from the canonical layout.
-A GPU-based sort generates an index list $\pi$, allowing the renderer to access parameters $\theta_{\pi(i)}$ while they remain stationary in memory.
-During backpropagation, gradients are scattered back to the canonical buffers:
+Maintaining optimizer consistency during depth-sorted rendering typically requires physically reordering parameter buffers, which incurs heavy data movement and can misalign optimizer states.
+We resolve this by decoupling the sorted view used for rendering/backprop from the canonical layout used for parameter storage and optimizer moments.
+A GPU-based sort produces a permutation $\pi$ that maps each sorted index $i$ to its canonical index $\pi(i)$.
+The renderer and backward kernels operate on the sorted view, while gradients are scattered back to canonical buffers:
 $\nabla\theta_{\pi(i)}\;{+}{=}\;g_{i},$ (10)
-ensuring that Adam optimizer states remain aligned with their respective parameters without CPU intervention or redundant memory copies.
-III-C3 On-GPU Parameter Optimization
-To eliminate the synchronization latency of CPU-driven updates, the entire optimization loop is executed on-device.
-We maintain Adam states (moments $m,v$) in dedicated GPU buffers and perform updates via a unified compute kernel.
-To ensure convergence under limited FP16 precision, we employ specialized parameterization: Opacity uses logit-space updates with a micro-step threshold; Scale is optimized in log-space with explicit clamping; and Rotation is updated via tangent-space projection and sign-unification.
-This fully GPU-resident schedule minimizes the host-device communication footprint while maximizing hardware utilization.
+ensuring that Adam optimizer states (stored in canonical order) remain aligned with their respective parameters, without CPU intervention or redundant memory copies.
 IV Experiments
@@
 Best results are in bold.
-Figure 4: Qualitative comparison on the LLFF dataset.
-Our method produces much sharper textures and more accurate thin structures (e.g., the leaves in Fern and the petals in Flower), closely matching the ground truth (GT).
-Figure 5: Qualitative comparison on the NeRF Synthetic dataset. In these object-centric synthetic scenes, our method produces high-fidelity results.
-Figure 6: Qualitative results on a mobile device across diverse datasets.
-We showcase real-time rendering screenshots from an iPhone 15, covering our self-collected scenes (rows 1-2), LLFF (row 3), and NeRF Synthetic (row 4).
-PocketGS consistently achieves high-fidelity reconstruction and sharp details across varying scene scales and capture conditions.
-The red labels indicate optimized point counts, demonstrating our method’s representation efficiency and robust generalization for practical on-device deployment.
 IV-A Experimental Setup
@@
 Finally, we report Count (the number of Gaussians) to characterize the trade-off between representation capacity and computational efficiency.
+Figure 4: Qualitative comparison on the LLFF dataset.
+Our method produces much sharper textures and more accurate thin structures (e.g., the leaves in Fern and the petals in Flower), closely matching the ground truth (GT).
 IV-B Main Results: Quality and Efficiency Comparison
@@
 PocketGS maintains high perceptual quality under mobile memory and bandwidth constraints by avoiding unnecessary densification while preserving critical structures.
+Figure 5: Qualitative comparison on the NeRF Synthetic dataset. In these object-centric synthetic scenes, our method produces high-fidelity results.
+Figure 6: Qualitative results on a mobile device across diverse datasets.
+We showcase real-time rendering screenshots from an iPhone 15, covering our self-collected scenes (rows 1-2), LLFF (row 3), and NeRF Synthetic (row 4).
+PocketGS consistently achieves high-fidelity reconstruction and sharp details across varying scene scales and capture conditions.
+The red labels indicate optimized point counts, demonstrating our method’s representation efficiency and robust generalization for practical on-device deployment.
 IV-C Memory Footprint
@@
 V Conclusion
-We introduce PocketGS, a fully on-device 3DGS training paradigm that co-designs geometry priors ($\mathcal{G}$), prior-conditioned initialization ($\mathcal{I}$), and hardware-aligned splatting optimization ($\mathcal{T}$) to make high-fidelity 3D reconstruction practical under mobile compute and memory constraints.
-PocketGS achieves perceptual quality competitive with or superior to workstation baselines while meeting smartphone budgets (e.g., $\sim$4-minute end-to-end training and <3<3 GB peak memory on iPhone 15), enabling practical professional-grade mobile 3D content creation for mixed reality and digital twins.
+In this paper, we present PocketGS, a fully on-device training framework for 3D Gaussian Splatting that targets the practical constraints of mobile devices.
+PocketGS is built upon three co-designed operators: (G) geometry-prior construction to provide compact and reliable geometric guidance, (I) prior-conditioned Gaussian parameterization to improve initialization and optimization under iteration-limited budgets, and (T) hardware-aligned splatting optimization to ensure efficient and stable differentiable splatting on mobile GPUs.
+Extensive experiments on public benchmarks and our MobileScan dataset show that PocketGS achieves strong perceptual quality while meeting smartphone budgets, e.g., approximately 4-minute end-to-end training with <3<3 GB peak memory on iPhone 15.
+These results indicate that high-fidelity 3D reconstruction and mobile 3D content creation can be made practical on commodity devices.
 Appendix A Implementation Details